DanielleLove:

2x+6x^2 How is that equal 2x(1+3x)

Vocaloid:

factor out the original expression 2x and 6x^2 have 2x as a common factor 2x/2x = 1 6x^2/(2x) = 3x therefore 2x+6x^2 = 2x(1+3x)

Vocaloid:

if you are still in doubt try applying the distributive property to 2x(1+3x)

DanielleLove:

ohhh ok I got it now thank you!

ILoveXXXTENTACION:

i love your profile pic, its so cute!!!